# Parameterized Controls and Isomorphic Representations of G-Expectations

### v1

@@ Abstract
This paper establishes a fundamental isomorphism between the abstract framework of $G$-expectation and a concrete parameterized family of classical stochastic processes, providing a unified framework for nonlinear stochastic calculus under model uncertainty. We demonstrate that $G$-Brownian motion and related stochastic differential equations (SDEs) can be intrinsically represented as controlled processes $\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$, where $\theta$ denotes an admissible control process taking values in a compact convex parameter set $\Theta$ encoding volatility uncertainty.
The core contribution is the construction of a complete algebraic and analytic isomorphism $\Phi$ that preserves all operations including addition, multiplication, function composition, and limit transitions. This isomorphism provides a rigorous mathematical foundation for understanding $G$-martingales, with a distinction between symmetric martingales (fair in all volatility scenarios) and asymmetric martingales (favorable but fair only in the worst-case scenario).
We develop a comprehensive theory for solving $G$-SDEs through their parameterized representations, establishing existence and uniqueness results under standard Lipschitz and growth conditions. The framework reveals how bounded control sets preserve these classical conditions while capturing volatility uncertainty.
The parameterized control perspective offers both mathematical rigor and computational practicality, enabling the transfer of classical stochastic calculus results to the $G$-expectation framework while maintaining the essential sublinear structure of model uncertainty.
Key Words: G-Expectation, G-Brownian Motion, Parameterized Controls, Isomorphic Representations, Volatility Uncertainty, Dynamic Programming Principle
@@ 1 Introduction
The quantification and management of model uncertainty—often referred to as Knightian uncertainty or ambiguity—represents a fundamental challenge across several disciplines, most notably in mathematical finance and decision theory. Traditional stochastic calculus, built upon the foundation of a single, precisely known probability measure, is inherently limited in its capacity to address situations where the model itself is imperfectly known. This limitation is particularly acute in financial markets, where volatility—the key parameter in option pricing and risk management—is notoriously difficult to estimate and is subject to significant time-varying uncertainty.
The groundbreaking framework of $G$-expectation, pioneered by Peng , provides a rigorous mathematical foundation for modeling such uncertainty. By replacing the classical linear expectation with a sublinear operator, the $G$-framework successfully captures the essence of volatility ambiguity. Within this framework, the $G$-Brownian motion $B_{t}$ and related stochastic differential equations ($G$-SDEs) have been defined, with their quadratic variation $\langle B\rangle_{t}$ evolving within a prescribed interval $\underline{\sigma}^{2}t,\overline{\sigma}^{2}t]$, naturally encoding the uncertainty in volatility. This leads to a rich theory of nonlinear stochastic calculus .
A profound insight emerging from this framework is the intrinsic connection between $G$-expectations and stochastic control theory. Specifically, the $G$-expectation can be naturally represented as the value function of a stochastic control problem:
$\hat{\mathbb{E}}[\varphi(B_{T})]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(\int_{0}^{T}\theta_{s}dW_{s}\right)\right],$
where the control process $\theta_{s}$ takes values in a compact set $\Theta$ representing the volatility uncertainty. This representation naturally leads to a Hamilton-Jacobi-Bellman (HJB) equation through the Dynamic Programming Principle (DPP):
$\partial_{t}v+G(D_{x}^{2}v)=0,\quad v(T,x)=\varphi(x),$
where the nonlinearity $G(A)=\frac{1}{2}\sup_{\gamma\in\Theta}\text{tr}(A\gamma\gamma^{\top})$ encodes the uncertainty set. The DPP manifests infinitesimally as:
$v(t,x)=\sup_{\theta\in\mathcal{A}_{t,t+h}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[v\left(t+h,x+\int_{t}^{t+h}\theta_{s}dW_{s}\right)\right].$
Despite this elegant connection, the abstract nature of the $G$-expectation can present a barrier to intuitive understanding and practical application. A central question thus arises:
Can we establish a more explicit and operational connection between the abstract $G$-expectation space and the concrete framework of parameterized controlled processes?
This paper provides a definitive affirmative answer by constructing a complete isomorphism between these two domains. We demonstrate that the entire $G$-expectation framework admits a rigorous correspondence to a family of classical stochastic processes controlled by volatility parameters. The core idea is to represent a $G$-Brownian motion path as a family of classical Itô processes:
$\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}},\quad\text{where}\quad B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}.$
The sublinear $G$-expectation is then revealed to be the supremum over all these classical scenarios, providing a stochastic control interpretation that complements the PDE perspective.
This isomorphic correspondence, denoted by $\Phi$, is shown to preserve the entire algebraic and analytic structure—including addition, multiplication, function composition, and limit transitions—thereby bridging the abstract nonlinear theory with the well-understood classical theory.
Contributions and Outline: The main contributions of this paper are as follows:
  • We construct a precise isomorphism $\Phi$ between the $G$-expectation space and a space of parameterized controlled processes, providing a concrete representation that complements the HJB equation characterization (Section 3).
  • We develop a comprehensive theory for $G$-SDEs, proving existence and uniqueness of solutions through their parameterized representations, thereby connecting the SDE solutions with the dynamic programming framework (Section 4).
  • Through this isomorphism, we characterize the fundamental structure of $G$-martingales, introducing a distinction between symmetric martingales (fair in all volatility scenarios) and asymmetric martingales (favorable but fair only in the worst-case scenario) (Sections 5 .
  • We resolve subtle definability difficulties associated with conditional $G$-expectations, clarifying the interplay between supremum operations and conditional expectations under model uncertainty.
The paper is organized as follows: Section 2 lays the theoretical foundation, formalizing $G$-expectation as a value function and deriving its connection to dynamic programming and nonlinear PDEs. Section 3 introduces the fundamental isomorphism, the central concept of this work. Sections 4 and 5 delve into the solutions of $G$-SDEs and the properties of $G$-martingales from this new perspective, respectively. Finally, Section 6 concludes.
Our work not only deepens the theoretical understanding of the $G$-framework but also unlocks new computational pathways. By translating complex nonlinear problems into families of more tractable classical ones, this parameterized control perspective offers a powerful tool for both theoretical analysis and practical computation in the presence of model uncertainty, while maintaining a natural connection to the HJB and dynamic programming framework.
2 Theoretical Foundation
@@ 2.1 G-Expectation as a Value Function
The fundamental insight is that G-expectation can be represented as the value function of a stochastic control problem. For a G-Brownian motion $B_{t}$ and terminal payoff $\varphi\in C_{b}(\mathbb{R}^{d})$, we have:
@@ 2.2 Dynamic Programming and Nonlinear PDEs
3 The Fundamental Isomorphism: G-Brownian motion as Parameterized Controlled Processes
The central insight of this work is establishing a natural isomorphism between elements in the G-expectation space and parameterized families of controlled processes. Specifically:
@@ 3 The Fundamental Isomorphism: G-Brownian motion as Parameterized Controlled Processes
    The isomorphism: There exists a bijective correspondence:
@@
This isomorphism reveals that:
  • G-processes are intrinsically controlled trajectories parameterized by volatility uncertainty
  • The G-expectation framework provides a geometric encoding of model uncertainty
  • Solutions to G-SDEs emerge as universal families of classical SDE solutions
@@ 3.1.1 G-Normal Distribution
The cornerstone of the theory is the G-normal distribution . Let $\xi$ be a random variable. It is said to follow a G-normal distribution, denoted $\xi\sim\mathcal{N}(0,[\underline{\sigma}^{2},\overline{\sigma}^{2}])$, if for any $\phi\in C_{b}(\mathbb{R})$, its sublinear expectation is defined by the following nonlinear heat equation:
@@ 3.1.2 G-Brownian Motion
3.2 A Generalized Construction via Random Variables
This generator governs the evolution of expectations under the G-Brownian motion. The drift term is zero, reflecting the martingale property of $B_{t}$, and the diffusion term is given by the supremum over constant parameters. This establishes the foundational link: classical G-Brownian motion corresponds to a constant diffusion coefficient uncertainty set $\Theta=[\underline{\sigma},\overline{\sigma}]$.
@@ 3.2 A Generalized Construction via Random Variables
Let $(\Omega,\mathcal{F},P)$ be a probability space sufficient to support a standard Brownian motion $\{W_{t}\}_{t\geq 0}$. For a given compact, convex parameter set $\Lambda\subset\mathbb{R}^{n}$, we define two fundamental random objects:
1. Max-Distribution: For a parameter $\mu\in\Lambda$, define the random variable $\eta(\mu)=\mu$. The associated nonlinear expectation $\mathbb{E}[\eta]$ is defined for a test function $\phi$ as:
@@
2. G-Normal Distribution: For a parameter $\sigma\in\Lambda$, define the random variable $X(\sigma)=\sigma W_{1}$ (which is classically normal distributed, $X(\sigma)\sim N(0,\sigma^{2})$). The associated nonlinear expectation $\mathbb{E}[X]$ is defined as:
$\mathbb{E}[\phi(X)]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(X(\sigma))]=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{1})].$ (3.7)
This constitutes a G-normal distribution with the uncertainty set $\Lambda$.
3. Quasi G-Brownian Motion: We define a process $\{B_{t}\}_{t\geq 0}$ by its action on test functions. For any $\phi\in C_{l,Lip}(\mathbb{R})$, its nonlinear expectation is given by:
$\mathbb{E}[\phi(B_{t})]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t})].$ (3.8)
Furthermore, for a process of the form $H_{t}=bB_{t}+mt$ (where the drift $m$ is also subject to uncertainty within a convex set), the expectation is defined as:
$\mathbb{E}[\phi(H_{t})]:=\sup_{(b,m)\in\Lambda^{\prime}}E_{P}[\phi(bW_{t}+mt)],$ (3.9)
where $\Lambda^{\prime}$ is a compact convex set in $\mathbb{R}^{2}$.
These constructions are defined on a filtered probability space represented by the quadruple:
$(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{E})$ (3.10)
where:
  • $\Omega$ is the sample space, representing the set of all possible outcomes;
  • $\mathcal{F}$ is a $\sigma$-algebra on $\Omega$, representing the set of all measurable events;
  • $(\mathcal{F}_{t})_{t\geq 0}$ is a filtration, i.e., an increasing family of sub-$\sigma$-algebras of $\mathcal{F}$ satisfying $\mathcal{F}_{s}\subseteq\mathcal{F}_{t}$ for all $0\leq s\leq t$, which represents the accumulation of information over time;
  • $\mathbb{E}$ is the sublinear expectation operator defined in equations (3.6) and (3.7), which generalizes the classical linear expectation to account for model uncertainty.
Remark 3.1.
The construction of the ”Quasi G-Brownian Motion” and the related objects herein represents the core methodological approach of this paper. It is intended to schematically illustrate the central idea of handling model uncertainty via the supremum over linear expectations. It is important to note that the process $\{B_{t}\}$ defined here is for methodological demonstration purposes; while it shares the same one-dimensional distributions with the strict mathematical G-Brownian motion, their joint distributions are entirely different, and thus they are not mathematically equivalent.
Definition 3.2 (General Construction of Sublinear Expectation via Model Uncertainty).
Let $(\Omega,\mathcal{F},P)$ be a classical probability space. Let $\Theta$ be a non-empty, compact, and convex parameter set representing the domain of model uncertainty.
Consider a generalparameter-controlled random variable $X(\theta)$, where $\theta\in\Theta$ is the parameter that dictates the model specification. The associated sublinear expectation $\mathbb{E}[\cdot]$ is defined for a test function $\phi$ (typically in the space of Lipschitz continuous functions $C_{l,Lip}(\mathbb{R}^{d})$) by the following operation:
$\mathbb{E}[\phi(X)]:=\sup_{\theta\in\Theta}E_{P}[\phi(X(\theta))].$ (3.11)
This means the operator $\mathbb{E}$ acts on the random variable $X$ by a two-step procedure:
1. For a fixed parameter $\theta\in\Theta$, compute the classical linear expectation $E_{P}$ of the random variable $\phi(X(\theta))$ under the probability measure $P$.
2. Subsequently, take the supremum of these resulting real numbers over all admissible parameters $\theta$ in the uncertainty set $\Theta$.
This operator is defined on the sublinear expectation space:
$(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{E})$ (3.12)
where:
  • $\Omega$ is the sample space;
  • $\mathcal{F}$ is a $\sigma$-algebra on $\Omega$;
  • $(\mathcal{F}_{t})_{t\geq 0}$ is a filtration representing the information flow over time;
  • $\mathbb{E}$ is the sublinear expectation operator defined above.
Remarks:
  • This general definition subsumes the specific constructions (e.g., Max-Distribution, G-Normal Distribution, Generalized G-Brownian Motion) presented in Definition 3.1 as special cases by choosing the parameter-controlled variable $X(\theta)$ and the uncertainty set $\Theta$ appropriately.
  • The operator $\mathbb{E}$ defined in this way is sublinear: $\mathbb{E}[X+Y]\leq\mathbb{E}[X]+\mathbb{E}[Y]$ and $\mathbb{E}[\lambda X]=\lambda\mathbb{E}[X]$ for $\lambda\geq 0$. It also preserves constants: $\mathbb{E}[c]=c$ for any constant $c$.
While the parameterized construction provides valuable insight, it is crucial to recognize its fundamental limitation: constant control sets cannot capture the path-dependent nature of true G-Brownian motion.
Consider the parameterized ”quasi G-Brownian motion” defined by:
$B_{t}^{\sigma}=\sigma W_{t},\quad\sigma\in\Lambda$ (3.13)
with nonlinear expectation:
$\mathbb{E}[\phi(B_{t})]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t})]$ (3.14)
Key observations:
  1. 1.
    One-dimensional distributions match: For fixed $t$,
    $\text{Law}_{G}(B_{t})\cong\sup_{\sigma\in\Lambda}\text{Law}_{P}(\sigma W_{t})$
    meaning marginal distributions coincide with true G-Brownian motion.
  2. 2.
    Path structure diverges: The joint distributions differ fundamentally:
    $\mathbb{E}[\phi(B_{t},B_{s})]=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t},\sigma W_{s})]\neq\hat{\mathbb{E}}[\phi(B_{t},B_{s})]$
    This occurs because constant $\sigma$ cannot capture the time-consistent volatility uncertainty of true G-Brownian motion.
  3. 3.
    The Difficulty in Defining Conditional Expectation
3.3 True G-Brownian
To overcome these limitations, we must introduce time-varying adaptive controls:
Definition 3.3 (True G-Brownian Motion via Adaptive Controls).
Let $\Theta\subset\mathbb{R}^{d\times d}$ be compact convex. Define:
@@ 3.3 True G-Brownian
$d\langle B^{\theta}\rangle_{t}=(\theta_{t}\theta_{t}^{\top})dt$ (3.16)
@@
For the G-Brownian motion $B_{t}$, the conditional expectation is:
${\mathbb{E}}_{t}[\varphi(B_{T})]=\sup_{\theta\in\mathcal{A}_{t,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(\int_{t}^{T}\theta_{s}dW_{s}\right)\middle|\mathcal{F}_{t}\right]$
This definition provides the mathematical foundation for conditional expectations in the G-framework, capturing the essence of model uncertainty through the supremum over volatility controls.
@@
For a G-Brownian motion $\{B_{t}\}_{t\geq 0}$ and a test function $\varphi\in C_{b.Lip}(\mathbb{R}^{d})$, the conditional G-expectation operator $\hat{\mathbb{E}}_{s}[\cdot]$ is defined as:
${\mathbb{E}}_{s}[\varphi(B_{t})]=\sup_{\theta\in\mathcal{A}_{s,t}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{s}+\int_{s}^{t}\theta_{u}dW_{u}\right)\middle|\mathcal{F}_{s}\right]$
@@
  • $\mathcal{A}_{s,t}^{\Theta}$ denotes admissible control processes on $[s,t]$
  • $\theta_{u}\in\Theta=[\underline{\sigma},\overline{\sigma}]$ represents volatility controls
  • $W_{u}$ is a standard Brownian motion under $\mathbb{P}$
@@
$\xi=\varphi(B_{t_{1}},B_{t_{2}}-B_{t_{1}},\cdots,B_{t_{n}}-B_{t_{n-1}}),\quad\varphi\in C_{b.Lip}(\mathbb{R}^{d\times n})$
@@
${\mathbb{E}}_{s}[\xi]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}}^{\theta},B_{t_{2}}^{\theta}-B_{t_{1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\middle|\mathcal{F}_{s}\right]$
where $B_{t}^{\theta}=\int_{0}^{t}\theta_{u}dW_{u}$ is the controlled process.
${\mathbb{E}}[\xi]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}}^{\theta},B_{t_{2}}^{\theta}-B_{t_{1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\right]$
@@
where for $t_{k-1}\leq s$$t_{k}>s$, we treat $B_{t_{i}}$ as constants $x_{i}$ in the expectation.
Note that our definition of conditional expectation here is completely consistent with the classical definition of G-conditional expectation under the definition of cylindrical Brownian motion, so we can similarly extend it to its completion space.

### v7

# A Parametric Methodology for $G$-Expectations and Stochastic Systems under Model Uncertainty

@@ Abstract
This paper develops a systematic parametric method for analyzing stochastic systems under volatility uncertainty within the $G$-expectation framework. Leveraging the dual representation of the $G$-expectation as a supremum over a family of probability measures, we operationalize the correspondence between the $G$-expectation space and a parameterized family of classical stochastic processes. Our approach decomposes complex nonlinear analysis into two distinct phases: a “linear implementation phase,” which utilizes classical stochastic analysis under a fixed parameter, and a “consistent estimation phase” across the parameter space.
We rigorously prove that this parametric mapping preserves fundamental stochastic structures, including the Itô integral and quadratic variation, thus enabling a direct transplantation of classical Itô calculus techniques into the $G$-framework. The advantages of this methodology are illustrated through its application to $G$-stochastic differential equations and $G$-backward stochastic differential equations. Furthermore, we demonstrate that this method effectively mitigates the inherent accumulation of estimation errors found in traditional stepwise sublinear procedures, yielding significantly more accurate estimates for fundamental quantities in robust stochastic analysis.
Keywords: $G$-Expectation, Parametric Representation, Model Uncertainty, $G$-SDEs, Uniform Estimation, Financial Applications, Precision Improvement
@@ 1 Introduction
The mathematical treatment of Knightian uncertainty, particularly model ambiguity in volatility, has been fundamentally transformed by the pioneering work of Peng and colleagues through the establishment of the rigorous framework of $G$-expectation theory [1, 3, 4]. This groundbreaking work provides a powerful, axiomatic foundation for robust financial analysis, enabling the pricing and hedging of derivatives, portfolio optimization, and risk assessment in environments where the underlying volatility is not precisely known but resides within a deterministic set $\underline{\sigma},\overline{\sigma}]$. The $G$-expectation framework offers a coherent apparatus with deep connections to stochastic control and fully nonlinear partial differential equations (PDEs), forming a cornerstone for modern quantitative finance under model uncertainty.
Furthermore, significant advances in related mathematical domains have enriched the landscape of stochastic analysis. The framework of quasi-sure analysis developed by Soner, Touzi, and Zhang [9] provides a powerful methodology for aggregating stochastic processes under a non-dominated family of measures, addressing fundamental questions of consistency and robustness. In a different direction, Malliavin calculus [42] has emerged as a profound tool for the fine analysis of probability laws on Wiener space, enabling deep sensitivity analysis and the study of regularity for solutions to stochastic systems.
With continuing research, the axiomatization and intrinsic approach to nonlinear expectation have witnessed significant advances. However, as research has matured and delved deeper into the intrinsic properties of the $G$-framework, a subtle yet significant divergence from the vast toolkit of classical stochastic analysis has emerged. While the foundational literature—notably Denis, Hu, and Peng [2]—clearly establishes the dual representation of the $G$-expectation as a supremum over a family of linear expectations, the systematic operationalization of this representation for stochastic calculus has remained largely underdeveloped. The pathwise dynamics of $G$-processes are often obscured when one works directly with sublinear evaluations or the associated fully nonlinear PDEs.
This gap has fostered a methodological conservatism in the field. There exists a justifiable, yet mathematically constraining, apprehension: the fear that stepping outside the intrinsic, nonlinear formalism might compromise the robustness of the analysis. This “mathematical fear” has, in practice, led to an over-reliance on stepwise sublinear procedures. While robust by construction, these methods often forcibly decouple the correlation structure between stochastic components by applying sublinear inequalities at intermediate steps. As we demonstrate in Section 7, this not only leads to a significant accumulation of estimation errors, yielding excessively conservative and distorted bounds, but also renders the analysis computationally inflexible and inaccessible to those well-versed in classical theory.
This paper directly addresses this core challenge by constructing a rigorous and systematic parametric methodology that bridges the $G$-expectation space with the classical probability space. Our primary contribution is not in identifying the duality—which is known—but in rigorously operationalizing it. We explicitly construct a homomorphism $\Phi$ that maps the $G$-expectation space to parameterized families of classical stochastic processes, and we prove that this mapping preserves the essential differential structures, including Itô integrals, quadratic variations, and the dynamics of stochastic differential equations.
Our framework strategically disentangles the analytical complexity by establishing a clear methodological dichotomy:
  1. 1.
    Linear Execution Phase: We demonstrate that the core computational workload—solving SDEs, applying Itô’s lemma, establishing pathwise properties—can be legitimately carried out within the familiar realm of classical linear stochastic analysis for each fixed parameter $\theta$. This allows researchers to leverage the entire arsenal of classical probability theory.
  2. 2.
    Uniform Estimation Phase: The crucial synthesis into the robust $G$-framework is achieved by a final, well-defined supremum operation, predicated on establishing estimates that hold uniformly across the parameter space $\Theta$.
This structured separation is transformative. It conclusively demonstrates that the “mathematical fear” of losing robustness is unfounded; robustness is guaranteed by the final supremum, provided the intermediate analysis is valid for each scenario. By consolidating the analytical challenge solely into the requirement of uniform estimates, our framework dramatically reduces complexity and, as empirically shown in Section 7, significantly improves estimation precision. It allows one to enjoy the robustness of the $G$-framework without sacrificing the power and flexibility of classical stochastic calculus.
The remainder of this paper is organized as follows. Section 2 revisits the foundational concepts and establishes the parametric representation. Section 3 is devoted to the rigorous construction of the homomorphism $\Phi$ and the proof of its structure-preserving properties. Sections 4 and 5 demonstrate the application of this methodology to $G$-SDEs and $G$-BSDEs, respectively. Section 6 provides a unified perspective on the homomorphism as a change of probabilistic lenses on the canonical path space. Section 7 presents a comparative analysis that quantitatively demonstrates the precision improvement of our approach. Section 8 discusses the broader implications of this parametric bridge for future research, and Section 9 concludes.
2 Fundamental Concepts and Parametric Representation
@@ 2.1 G-Expectation as a Value Function
The fundamental insight is that G-expectation can be represented as the value function of a stochastic control problem[1]. For a G-Brownian motion $B_{t}$ and terminal payoff $\varphi\in C_{b}(\mathbb{R}^{d})$, we have:
@@ 2.2 Dynamic Programming and Nonlinear PDEs
The central insight of this work is establishing a natural homomorphism between elements in the G-expectation space and parameterized families of controlled processes. Specifically:
@@
    The homomorphism: There exists a injective correspondence:
@@ 2.3.1 G-Normal Distribution
The cornerstone of the theory is the G-normal distribution . Hereafter, we will focus on the one-dimensional case, while the analysis for multi-dimensional situations follows similarly.Let $\xi$ be a random variable. It is said to follow a G-normal distribution, denoted $\xi\sim\mathcal{N}(0,[\underline{\sigma}^{2},\overline{\sigma}^{2}])$, if for any $\phi\in C_{b}(\mathbb{R})$, its sublinear expectation is defined by the following nonlinear heat equation:
@@ 2.3.2 G-Brownian Motion
2.4 Parameterized Control Representation
Now we draw upon Peng’s framework to construct the corresponding parameterized control representation.For details of the construction, see [1].
@@ 2.4 Parameterized Control Representation
Let $(\Omega,\mathcal{F},\mathcal{F}_{t},P)$ be a filtered probability space sufficient to support a standard Brownian motion $\{W_{t}\}_{t\geq 0}$. For a given compact, convex parameter set $\Lambda\subset\mathbb{R}^{n}$ and set $\Theta\subset\mathbb{R}^{d\times d}$, we define two fundamental random objects:
Max-Distribution: For a parameter $\mu\in\Lambda$, define the random variable $\eta(\mu)=\mu$. The associated nonlinear expectation $\mathbb{E}[\eta]$ is defined for a test function $\phi$ as:
@@
G-Brownian Motion : Let $\Theta\subset\mathbb{R}^{d\times d}$ be compact convex. Define:
@@
$d\langle B^{\theta}\rangle_{t}=\theta_{t}^{2}dt$ (2.13)
@@
Remark 2.1.
The compactness of the control set $\Theta$ ensures the existence of a countable dense subset $\Theta_{d}\subset\Theta$, which allows us to approximate the supremum over all controls through this subset:
$\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{P}\left[\varphi\left(\int_{0}^{T}\theta_{s}dW_{s}\right)\right]=\sup_{\theta\in\mathcal{A}^{\Theta_{d}}}\mathbb{E}_{P}\left[\varphi\left(\int_{0}^{T}\theta_{s}dW_{s}\right)\right].$
This approximation property stems from the $L^{2}$-continuous dependence on the control parameter and provides analytical convenience in various contexts. The introduction of such topological considerations within our parametric framework can facilitate certain analytical techniques, though this aspect is not central to our main development.
@@
For a G-Brownian motion $\{B_{t}\}_{t\geq 0}$ and a test function $\varphi\in C_{b,Lip}(\mathbb{R})$, the conditional G-expectation operator $\hat{\mathbb{E}}_{s}[\cdot]$ is defined as:
$\mathbb{E}_{s}[\varphi(B_{t})]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{s}+\int_{s}^{t}\theta_{u}dW_{u}\right)\Big|\mathcal{F}_{s}\right]$
@@
  • $\mathcal{A}_{s,T}^{\Theta}$ denotes admissible control processes on $[s,T]$,
  • $\theta_{u}\in\Theta=[\underline{\sigma},\overline{\sigma}]$ represents volatility controls,
  • $W_{u}$ is a standard Brownian motion under $\mathbb{P}$.
@@
$\xi=\varphi(B_{t_{1}},B_{t_{2}}-B_{t_{1}},\cdots,B_{t_{n}}-B_{t_{n-1}}),\quad\varphi\in C_{b,Lip}(\mathbb{R}^{n})$
@@
$\mathbb{E}_{s}[\xi]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}},B_{t_{2}}-B_{t_{1}},\cdots,B_{t_{n}}-B_{t_{n-1}}\right)\Big|\mathcal{F}_{s}\right]$
Note: In the above, the control process $\theta$ affects the increments after time $s$. For increments before time $s$, they are fixed by the observation.
@@
where for $t_{k-1}\leq s<t_{k}$, we treat $B_{t_{i}}$ for $i=1,\dots,k-1$ as constants $x_{i}$ in the expectation, and for $t\geq s$, we define $B_{t}^{\theta}=B_{s}+\int_{s}^{t}\theta_{u}dW_{u}$.
$\mathbb{E}[\xi]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}}^{\theta},B_{t_{2}}^{\theta}-B_{t_{1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\right].$
Note that our definition of conditional expectation is completely consistent with the classical definition of G-conditional expectation under cylindrical Brownian motion, allowing for natural extension to the completion space. This framework extends seamlessly to stopping times, where conditioning occurs on the information up to the stopping time $\tau$. Furthermore, the elements preceding the conditional expectation, as members of the G-expectation space, naturally correspond to elements in the parameterized family. For notational simplicity, we have omitted explicit parameterization in these expressions, though the underlying parametric structure remains intact throughout our framework.